# The T2K Collaboration

### v1

@@ I INTRODUCTION
In order to improve our understanding of the phenomenon of neutrino oscillations, a new generation of experiments is required with increased detector mass, higher-power beam and better near detectors yielding improved statistics and sensitivity. Within the paradigm of three Dirac neutrinos with their flavor states being linear combinations of their mass eigenstates as determined by the PMNS mixing matrix [1], the remaining major unknowns are: the ordering of the mass eigenstates and the value of the CP symmetry violating parameter. Together with the hypothetical existence of sterile neutrinos these are the main questions to be addressed by several long- and short-baseline neutrino oscillation experiments. The precision of measurements performed in experiments like T2K [2], NOvA [3], Hyper-Kamiokande (HK) [4] and DUNE [5] relies strongly on reducing the systematic uncertainties originating from our limited knowledge of neutrino-nucleus cross sections [6]. T2K uses a neutrino beam peaked in the sub-GeV energy range hence the main neutrino interaction channel is Charged Current Quasi-Elastic (CCQE) scattering with a nucleon. This interaction mode is described with the standard Llewellyn Smith formalism [7] but it is far from being fully understood due to the axial structure of the weak current and complications introduced through nuclear effects, as neutrino oscillation experiments use complex nuclear targets such as carbon, water, or argon. Other interaction modes bring their own unknowns and this motivates experimental efforts to measure neutrino cross sections with different fluxes, targets and definitions of the experimental signal.
The T2K experiment uses Super-Kamiokande, a water Cherenkov detector [8], as the far detector to measure neutrinos after oscillation and ND280 [9] as the near detector, measuring the neutrino flux before oscillation. The primary target material in ND280 is hydrocarbon, therefore it is highly desirable to also measure neutrino interactions on a water target at the near detector site. WAGASCI-BabyMIND with its mixed water and hydrocarbon composition allows the simultaneous measurement of the neutrino cross section on both of these target materials, motivating its construction and inclusion within the T2K experiment.
The WAGASCI-BabyMIND is the collective name of the detectors shown in Fig. 1. It was installed in 2018 at the main ND280 [10] complex in a position 1.5 degrees off-axis on the B2 floor of the Neutrino Monitoring (NM) building of the Japan-Proton Accelerator Research Complex (J-PARC). Together, the WAter Grid And SCIntillator modules (WAGASCIs), Proton Module (PM), Wall Muon Range Detectors (WMs) and Baby Magnetized Iron Neutrino Detector (BabyMIND) form an integrated system of detectors. In this paper, we present new cross section measurements made by the WAGASCI-BabyMIND on both H2O and CH targets.
@@
The results presented in this paper follow previous cross section measurements on a standalone WAGASCI module [11]. Here we report the first measurement with the full WAGASCI-BabyMIND set up. The results reported in the paper are the first obtained using a new near detector in T2K with a water component and measures the Carbon/Oxygen cross section ratio at a higher energy than ND280.
@@ II The T2K experiment
The T2K experiment has achieved several major milestones, including the first observation of electron neutrino appearance in a muon neutrino beam [12] and precision measurements of $\theta_{23}$—a key parameter in the PMNS mixing matrix [13]. Moreover, T2K recently demonstrated sensitivity to measuring $\delta_{CP}$, a parameter of the PMNS matrix responsible for the violation of the CP symmetry in neutrino interactions [14, 15]. These observations are crucial for understanding the fundamental properties of neutrinos and their role in the universe [16].
@@ II.1 The muon neutrino beam
The T2K experiment utilizes a muon neutrino beam produced at the J-PARC facility. Protons are accelerated to 30 GeV in stages by three accelerators [17]. These high-energy protons collide with a graphite target producing secondary particles, predominantly pions. The pions are focused in the forward direction by three magnetic horns, which also defocus particles with the wrong charge, thereby enhancing the purity of the neutrino beam. As these pions travel through a 96-meter-long decay pipe, they decay predominantly producing muon neutrinos or muon anti-neutrinos in either Forward Horn Current or Reverse Horn Current mode.
@@ II.2 The WAGASCI-BabyMIND
Near detectors, such as the WAGASCI-BabyMIND [11] shown in Fig. 1 and ND280 [10], measure the unoscillated neutrino flux. This allows for detailed measurements of neutrino interactions and cross sections before oscillation, helping to constrain the oscillation measurements at the far detector and deepening our understanding of neutrino interactions in the energy range of interest.
@@
The WAGASCIs serve as the primary water target for neutrinos and consists of a three-dimensional grid of scintillator immersed in water, allowing for high signal purity. The internal structure and dimensions are explained in detail in Ref.  [11].
@@
The fiducial mass of the PM is 313 kg of CH and that of the WAGASCIs is 229 kg of H2O and 62 kg of CH in total. By combining events from both the WAGASCIs and the PM in the cross section fit, we can decorrelate the cross sections on water and hydro-carbon. A similar strategy was employed by the ND280 detector as described in Ref. [18]. The geometry of the PM is briefly discussed in Ref. [2], while the performance of the scintillator bars is described in Ref. [19].
A preliminary cross section measurement [11] was performed between 2017 and 2018 using only one WAGASCI module and the PM as target detectors and an INGRID module [20] as a downstream muon range detector. In the measurement presented in this paper, we have improved the detectable phase space of the downrange muons by replacing the INGRID module with two new detectors:
@@
    One BabyMIND detector, installed downstream of the WAGASCIs and made of 33 magnetized iron planes and 18 scintillator planes. It provides the capability to identify the charge of muons by the curvature of their tracks and precisely measure their momentum by measuring their penetration distance. A more in-depth description of the structure and performance can be found in [21, 22]. The BabyMIND and WAGASCIs are not aligned along the flux axis because of the narrow available space in the B2 pit.
@@ III.2 Monte-Carlo simulation
The selection criteria were studied using Monte Carlo (MC) simulations. A MC framework was developed to generate simulated datasets for this analysis with the approach consistent with that adopted in the previous analysis [11]. Simulated datasets for each interaction target were produced and included the initial neutrino flux properties, particle information, hit/track information, etc. The T2K neutrino flux at 1.5 degrees off-axis was simulated using the T2K package JNUBEAM (version 2021.2) [23]. The output of the simulated flux information was fed into the T2K neutrino interaction generator, NEUT (version 5.3.2) [24]. Using the kinematic variables of the generated final state particles of the simulated neutrino-nucleus interaction as input, the detector response was then simulated by Geant4 (version 4.10.6.2) [25] with the ‘QGSP BERT’ model which also simulates the subsequent secondary or tertiary interactions within the detector material.
@@ III.3 Reconstruction
The reconstruction techniques were based on those implemented for the previous analysis [11] but extended so that the methods could be applied to all the detectors including the magnetized detector, BabyMIND. The modifications were done by tuning the reconstruction parameters on a detector-by-detector basis. The track reconstruction process was composed of three steps: track seeding, track matching and reconstruction of track properties. The track seeding tries to find track candidates using hit information. After track candidates were found in each detector, tracks were connected between multiple detectors. Once a matched track was reconstructed, the algorithm reconstructs a vertex by identifying the most upstream hit of the matched track. Finally, we extracted track parameters such as momentum, angle, charge and particle type based on the reconstructed tracks.
@@ III.9 Number of target nucleons
where $j$ runs over each reconstructed bin for all samples and $N^{\mathrm{exp}}(N^{\mathrm{obs}})$ is the number of selected events predicted by MC (obtained in the data), $\beta_{j}$ are the Barlow-Beeston scaling parameters which account for the uncertainty of finite MC simulation (for the definition see [28]) and $\sigma$ is the MC statistical uncertainty.
@@ V.1 Neutrino Flux Systematics
Neutrino flux uncertainties were evaluated using a method employed previously in other T2K studies [23]. The covariance matrix binning has been chosen to ensure approximately equal statistics in each bin as can be seen in Tab. 2 and the resultant matrix is shown in Fig. 10.
@@
In Fig. 11 the fractional error on the muon neutrino flux is shown by the colored spectrum. In the background, the curve filled with gray shows the neutrino flux as a function of neutrino energy at the 1.5 off-axis angle. The uncertainties stem from several factors, including the hadron production model, the profile of the proton beam, the off-axis angle, the horn current, horn alignment, wrong-sign ($\nu/\bar{\nu}$) neutrino contamination and others. The T2K replica target hadron production measurements in NA61/SHINE [29, 30, 31] were used to define the hadron uncertainties reported here. How each source of error impacts the flux prediction is detailed in [23].
@@ V.2 Neutrino Interaction Systematics
Our MC simulation generated particles from interaction vertices based on theoretical models that incorporate our current understanding of neutrino interactions. However, these models rely on approximations due to limited experimental data and/or the complexity of underlying theoretical models. We used the NEUT software [24] to simulate neutrino interactions.
@@
Dial Name Prior Uncertainty Description
@@
MaCCQE 1.21 GeV -0.3, +0.3 CCQE axial mass
QETwk_High_Q2_Weight_1 1 -0.11, +0.11 Reweight CCQE events with $0.25<Q^{2}<0.5$
QETwk_High_Q2_Weight_2 1 -0.18, +0.18 Reweight CCQE events with $0.5<Q^{2}<1$
QETwk_High_Q2_Weight_3 1 -0.4, +0.4 Reweight CCQE events with $Q^{2}>1$
SF_OptPotTwkDial_O16 0 -0, +0.49 SF Optical Potential Correction, O
SF_OptPotTwkDial_C12 0 -0, +0.49 SF Optical Potential Correction, C
SF_P1_2Shell_MeanF_Norm_O 0 -0.2, +0.2 Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_P1_2Shell_MeanF_PMissShape_O 0 -1, +1 Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
SF_P3_2Shell_MeanF_Norm_O 0 -0.4, +0.4 Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_P3_2Shell_MeanF_PMissShape_O 0 -1, +1 Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
SF_PBTwkDial_Hybrid_C12_nu 0 -2, +0.6 Hybrid nucl. mom cut off and template normalisation in $q_{3}$, $q_{0}$
SF_PBTwkDial_Hybrid_C12_nubar 0 -2, +0.6 Hybrid nucl. mom cut off and template normalisation in $q_{3}$, $q_{0}$
SF_PBTwkDial_Hybrid_O16_nu 0 -2, +0.6 Hybrid nucl. mom cut off and template normalisation in $q_{3}$, $q_{0}$
SF_PBTwkDial_Hybrid_O16_nubar 0 -2, +0.6 Hybrid nucl. mom cut off and template normalisation in $q_{3}$, $q_{0}$
SF_PShell_MeanF_Norm_C 0 -0.2, +0.2 Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_PShell_MeanF_PMissShape_C 0 -1, +1 Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
SF_SRC_Norm_C 1 -1, +1 Overall normalisation in some range of $p_{m}$, $E_{m}$
SF_SRC_Norm_O 1 -1, +1 Overall normalisation in some range of $p_{m}$, $E_{m}$
SF_SShell_MeanF_Norm_C 0 -0.4, +0.4 Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_SShell_MeanF_Norm_O 0 -0.2, +0.2 Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_SShell_MeanF_PMissShape_C 0 -1, +1 Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
SF_SShell_MeanF_PMissShape_O 0 -1, +1 Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
@@
MECTwkDial_Norm_C12 1 -0.99, +0.99 2p2h Normalization Carbon
MECTwkDial_Norm_O16 1 -0.99, +0.99 2p2h Normalization Oxygen
MECTwkDial_Norm_Other 1 -0.49, +0.49 2p2h Normalization Others
MECTwkDial_PDDWeight_C12_NN 0 +0.49, -0.49 2p2h shape on Carbon: contributions change for NN pairs
MECTwkDial_PDDWeight_C12_np 0 -0.49, +0.49 2p2h shape on Carbon: contributions change for NP pairs
MECTwkDial_PDDWeight_O16_NN 0 -0.49, +0.49 2p2h shape on Oxygen: contributions change for NN pairs
MECTwkDial_PDDWeight_O16_np 0 -0.49, +0.49 2p2h shape on Oxygen: contributions change for NP pairs
MECTwkDial_PNNN_Shape 0 -0.33, +0.33 2p2h shape: NN or Np nucleon pair
@@
RES_E_b_C_numu 25 MeV -24.9, +24.9 Resonant SPP $E_{b}$ value (MeV) for muon neutrinos on Carbon
RES_E_b_O_numu 25 MeV -24.9, +24.9 Resonant SPP $E_{b}$ value (MeV) for muon neutrinos on Oxygen
RES_E_b_C_numubar 25 MeV -24.9, +24.9 Resonant SPP $E_{b}$ value (MeV) for $\overline{\nu}_{\mu}$ on Carbon
RES_E_b_O_numubar 25 MeV -24.9, +24.9 Resonant SPP $E_{b}$ value (MeV) for $\overline{\nu}_{\mu}$ on Oxygen
BgSclRES 1.3 -0.15, +0.15 I½ non-resonant background
CA5RES 1.01 -0.15, +0.15 CA5(0): value at $Q^{2}=0$ of the axial form factor
MaRES 0.95 GeV -0.15, +0.15 RES axial mass
@@
PionFSI_AbsProb 1.404 -0.432, +0.432 Pion FSI absorption (FEFABS)
PionFSI_CExHighMomProb 1.8 -0.288, +0.288 Pion FSI Single charge exchange (high energy) (FEFCXH)
PionFSI_CExLowMomProb 0.697 -0.305, +0.305 Pion FSI Single charge exchange (low energy) (FEFCX)
PionFSI_InelProb 1.002 -1.101, +1.101 Pion FSI Hadron (N+n pi) production (FEFINEL)
PionFSI_QEHighMomProb 1.824 -0.859, +0.859 Pion FSI QE scattering (high energy) (FEFQEH)
PionFSI_QELowMomProb 1.069 -0.313, +0.313 Pion FSI QE scattering (low energy) (FEFQE)
TwkDial_FateNucleonFSI 0 -0.3, +0.3 Fate of final state nucleons after the FSI interactions
@@
CC_DIS_norm_nu 1 -0.15, +0.15 Normalisation of CC DIS for neutrinos
CC_DIS_norm_nubar 1 -0.15, +0.15 Normalisation of CC DIS for anti-neutrinos
CC_MultiPi_norm_nu 1 -0.2, +0.2 Multi-pion production normalisation (nu)
CC_MultiPi_norm_nubar 1 -0.2, +0.2 Multi-pion production normalisation (nubar)
Table 3: List of all systematic parameters from neutrino interaction models that were included in the analysis. The column ‘Dial Name’ is the name of the parameter as it appears in the source code. The column ‘Prior’ is the parameter’s prior value before variation. The column ‘Uncertainty’ represents the $\pm 1\sigma$ prior uncertainty. The column ‘Description’ contains a short description of the meaning of the parameter.

### v2

@@ I INTRODUCTION
In order to improve our understanding of the phenomenon of neutrino oscillations, a new generation of experiments is required with increased detector mass, higher-power beam and better near detectors yielding improved statistics and sensitivity. Within the paradigm of three Dirac neutrinos with their flavor states being linear combinations of their mass eigenstates as determined by the PMNS mixing matrix [PMNS], the remaining major unknowns are: the ordering of the mass eigenstates and the value of the CP symmetry violating parameter. Together with the hypothetical existence of sterile neutrinos these are the main questions to be addressed by several long- and short-baseline neutrino oscillation experiments. The precision of measurements performed in experiments like T2K [T2K:2011qtm], NOvA [Nova], Hyper-Kamiokande (HK) [HyperKamiokande] and DUNE [DUNE] relies strongly on reducing the systematic uncertainties originating from our limited knowledge of neutrino-nucleus cross sections [Avanzini:2021qlx]. T2K uses a neutrino beam peaked in the sub-GeV energy range hence the main neutrino interaction channel is Charged Current Quasi-Elastic (CCQE) scattering with a nucleon. This interaction mode is described with the standard Llewellyn Smith formalism [Llewellyn-Smith] but it is far from being fully understood due to the axial structure of the weak current and complications introduced through nuclear effects, as neutrino oscillation experiments use complex nuclear targets such as carbon, water, or argon. Other interaction modes bring their own unknowns and this motivates experimental efforts to measure neutrino cross sections with different fluxes, targets and definitions of the experimental signal.
The T2K experiment uses Super-Kamiokande, a water Cherenkov detector [SuperKamiokande], as the far detector to measure neutrinos after oscillation and ND280 [Kudenko:2008ia] as the near detector, measuring the neutrino flux before oscillation. The primary target material in ND280 is hydrocarbon, therefore it is highly desirable to also measure neutrino interactions on a water target at the near detector site. WAGASCI-BabyMIND with its mixed water and hydrocarbon composition allows the simultaneous measurement of the neutrino cross section on both of these target materials, motivating its construction and inclusion within the T2K experiment.
The WAGASCI-BabyMIND is the collective name of the detectors shown in Fig. 1. It was installed in 2018 at the main ND280 [T2K:2023qjb] complex in a position 1.5 degrees off-axis on the B2 floor of the Neutrino Monitoring (NM) building of the Japan-Proton Accelerator Research Complex (J-PARC). Together, the WAter Grid And SCIntillator modules (WAGASCIs), Proton Module (PM), Wall Muon Range Detectors (WMs) and Baby Magnetized Iron Neutrino Detector (BabyMIND) form an integrated system of detectors. In this paper, we present new cross section measurements made by the WAGASCI-BabyMIND on both H2O and CH targets.
@@
The results presented in this paper follow previous cross section measurements on a standalone WAGASCI module [10.1093/ptep/ptab014]. Here we report the first measurement with the full WAGASCI-BabyMIND set up. The results reported in the paper are the first obtained using a new near detector in T2K with a water component and measures the Carbon/Oxygen cross section ratio at a higher energy than ND280.
@@ II The T2K experiment
The T2K experiment has achieved several major milestones, including the first observation of electron neutrino appearance in a muon neutrino beam [T2K:2011ypd] and precision measurements of $\theta_{23}$—a key parameter in the PMNS mixing matrix [T2K:2014ghj]. Moreover, T2K recently demonstrated sensitivity to measuring $\delta_{CP}$, a parameter of the PMNS matrix responsible for the violation of the CP symmetry in neutrino interactions [Abe2020-Nature, PhysRevD.91.072010]. These observations are crucial for understanding the fundamental properties of neutrinos and their role in the universe [Canetti_2012].
@@ II.1 The muon neutrino beam
The T2K experiment utilizes a muon neutrino beam produced at the J-PARC facility. Protons are accelerated to 30 GeV in stages by three accelerators [10.1093/ptep/pts020]. These high-energy protons collide with a graphite target producing secondary particles, predominantly pions. The pions are focused in the forward direction by three magnetic horns, which also defocus particles with the wrong charge, thereby enhancing the purity of the neutrino beam. As these pions travel through a 96-meter-long decay pipe, they decay predominantly producing muon neutrinos or muon anti-neutrinos in either Forward Horn Current or Reverse Horn Current mode.
@@ II.2 The WAGASCI-BabyMIND
Near detectors, such as the WAGASCI-BabyMIND [10.1093/ptep/ptab014] shown in Fig. 1 and ND280 [T2K:2023qjb], measure the unoscillated neutrino flux. This allows for detailed measurements of neutrino interactions and cross sections before oscillation, helping to constrain the oscillation measurements at the far detector and deepening our understanding of neutrino interactions in the energy range of interest.
@@
The WAGASCIs serve as the primary water target for neutrinos and consists of a three-dimensional grid of scintillator immersed in water, allowing for high signal purity. The internal structure and dimensions are explained in detail in Ref.  [10.1093/ptep/ptab014].
@@
The fiducial mass of the PM is 313 kg of CH and that of the WAGASCIs is 229 kg of H2O and 62 kg of CH in total. By combining events from both the WAGASCIs and the PM in the cross section fit, we can decorrelate the cross sections on water and hydro-carbon. A similar strategy was employed by the ND280 detector as described in Ref. [PhysRevD.101.112004]. The geometry of the PM is briefly discussed in Ref. [T2K:2011qtm], while the performance of the scintillator bars is described in Ref. [NITTA2004147].
A preliminary cross section measurement [10.1093/ptep/ptab014] was performed between 2017 and 2018 using only one WAGASCI module and the PM as target detectors and an INGRID module [ABE2012211] as a downstream muon range detector. In the measurement presented in this paper, we have improved the detectable phase space of the downrange muons by replacing the INGRID module with two new detectors:
@@
    One BabyMIND detector, installed downstream of the WAGASCIs and made of 33 magnetized iron planes and 18 scintillator planes. It provides the capability to identify the charge of muons by the curvature of their tracks and precisely measure their momentum by measuring their penetration distance. A more in-depth description of the structure and performance can be found in [BabyMIND:2017mys, BabyMIND:2020lxx]. The BabyMIND and WAGASCIs are not aligned along the flux axis because of the narrow available space in the B2 pit.
@@ III.2 Monte-Carlo simulation
The selection criteria were studied using Monte Carlo (MC) simulations. A MC framework was developed to generate simulated datasets for this analysis with the approach consistent with that adopted in the previous analysis [10.1093/ptep/ptab014]. Simulated datasets for each interaction target were produced and included the initial neutrino flux properties, particle information, hit/track information, etc. The T2K neutrino flux at 1.5 degrees off-axis was simulated using the T2K package JNUBEAM (version 2021.2) [PhysRevD.87.012001]. The output of the simulated flux information was fed into the T2K neutrino interaction generator, NEUT (version 5.3.2) [Hayato:2002sd]. Using the kinematic variables of the generated final state particles of the simulated neutrino-nucleus interaction as input, the detector response was then simulated by Geant4 (version 4.10.6.2) [AGOSTINELLI2003250] with the ‘QGSP BERT’ model which also simulates the subsequent secondary or tertiary interactions within the detector material.
@@ III.3 Reconstruction
The reconstruction techniques were based on those implemented for the previous analysis [10.1093/ptep/ptab014] but extended so that the methods could be applied to all the detectors including the magnetized detector, BabyMIND. The modifications were done by tuning the reconstruction parameters on a detector-by-detector basis. The track reconstruction process was composed of three steps: track seeding, track matching and reconstruction of track properties. The track seeding tries to find track candidates using hit information. After track candidates were found in each detector, tracks were connected between multiple detectors. Once a matched track was reconstructed, the algorithm reconstructs a vertex by identifying the most upstream hit of the matched track. Finally, we extracted track parameters such as momentum, angle, charge and particle type based on the reconstructed tracks.
@@ III.9 Number of target nucleons
where $j$ runs over each reconstructed bin for all samples and $N^{\mathrm{exp}}(N^{\mathrm{obs}})$ is the number of selected events predicted by MC (obtained in the data), $\beta_{j}$ are the Barlow-Beeston scaling parameters which account for the uncertainty of finite MC simulation (for the definition see [PhysRevD.108.112009]) and $\sigma$ is the MC statistical uncertainty.
@@ V.1 Neutrino Flux Systematics
Neutrino flux uncertainties were evaluated using a method employed previously in other T2K studies [PhysRevD.87.012001]. The covariance matrix binning has been chosen to ensure approximately equal statistics in each bin as can be seen in Tab. 2 and the resultant matrix is shown in Fig. 10.
@@
In Fig. 11 the fractional error on the muon neutrino flux is shown by the colored spectrum. In the background, the curve filled with gray shows the neutrino flux as a function of neutrino energy at the 1.5 off-axis angle. The uncertainties stem from several factors, including the hadron production model, the profile of the proton beam, the off-axis angle, the horn current, horn alignment, wrong-sign ($\nu/\bar{\nu}$) neutrino contamination and others. The T2K replica target hadron production measurements in NA61/SHINE [Abgrall:2019EPJC.79.2, Abgrall:2016EPJC.76.11, Abgrall:2016EPJC.76.2] were used to define the hadron uncertainties reported here. How each source of error impacts the flux prediction is detailed in [PhysRevD.87.012001].
@@ V.2 Neutrino Interaction Systematics
Our MC simulation generated particles from interaction vertices based on theoretical models that incorporate our current understanding of neutrino interactions. However, these models rely on approximations due to limited experimental data and/or the complexity of underlying theoretical models. We used the NEUT software [Hayato:2002sd] to simulate neutrino interactions.
@@
Dial Name Prior Uncertainty Target Description
@@
MaCCQE 1.21 GeV -0.3, +0.3 Both CCQE axial mass
QETwk_High_Q2_Weight_1 1 -0.11, +0.11 Both Reweight CCQE events with $0.25<Q^{2}<0.5$
QETwk_High_Q2_Weight_2 1 -0.18, +0.18 Both Reweight CCQE events with $0.5<Q^{2}<1$
QETwk_High_Q2_Weight_3 1 -0.4, +0.4 Both Reweight CCQE events with $Q^{2}>1$
SF_OptPotTwkDial_O16 0 -0, +0.49 Water SF Optical Potential Correction, O
SF_OptPotTwkDial_C12 0 -0, +0.49 CH SF Optical Potential Correction, C
SF_P1_2Shell_MeanF_Norm_O 0 -0.2, +0.2 Water Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_P1_2Shell_MeanF_PMissShape_O 0 -1, +1 Water Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
SF_P3_2Shell_MeanF_Norm_O 0 -0.4, +0.4 Water Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_P3_2Shell_MeanF_PMissShape_O 0 -1, +1 Water Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
SF_PBTwkDial_Hybrid_C12_nu 0 -2, +0.6 CH Hybrid nucl. mom cut off and template normalisation in $q_{3}$, $q_{0}$
SF_PBTwkDial_Hybrid_C12_nubar 0 -2, +0.6 CH Hybrid nucl. mom cut off and template normalisation in $q_{3}$, $q_{0}$
SF_PBTwkDial_Hybrid_O16_nu 0 -2, +0.6 Water Hybrid nucl. mom cut off and template normalisation in $q_{3}$, $q_{0}$
SF_PBTwkDial_Hybrid_O16_nubar 0 -2, +0.6 Water Hybrid nucl. mom cut off and template normalisation in $q_{3}$, $q_{0}$
SF_PShell_MeanF_Norm_C 0 -0.2, +0.2 CH Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_PShell_MeanF_PMissShape_C 0 -1, +1 CH Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
SF_SRC_Norm_C 1 -1, +1 CH Overall normalisation in some range of $p_{m}$, $E_{m}$
SF_SRC_Norm_O 1 -1, +1 Water Overall normalisation in some range of $p_{m}$, $E_{m}$
SF_SShell_MeanF_Norm_C 0 -0.4, +0.4 CH Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_SShell_MeanF_Norm_O 0 -0.2, +0.2 Water Norm of Gaussian positioned at some $E_{m}$ for restricted $p_{m}$
SF_SShell_MeanF_PMissShape_C 0 -1, +1 CH Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
SF_SShell_MeanF_PMissShape_O 0 -1, +1 Water Template normalisation in $p_{m}$ in some range of $p_{m}$, $E_{m}$
@@
MECTwkDial_Norm_C12 1 -0.99, +0.99 CH 2p2h Normalization Carbon
MECTwkDial_Norm_O16 1 -0.99, +0.99 Water 2p2h Normalization Oxygen
MECTwkDial_Norm_Other 1 -0.49, +0.49 Both 2p2h Normalization Others
MECTwkDial_PDDWeight_C12_NN 0 +0.49, -0.49 CH 2p2h shape on Carbon: contributions change for NN pairs
MECTwkDial_PDDWeight_C12_np 0 -0.49, +0.49 CH 2p2h shape on Carbon: contributions change for NP pairs
MECTwkDial_PDDWeight_O16_NN 0 -0.49, +0.49 Water 2p2h shape on Oxygen: contributions change for NN pairs
MECTwkDial_PDDWeight_O16_np 0 -0.49, +0.49 Water 2p2h shape on Oxygen: contributions change for NP pairs
MECTwkDial_PNNN_Shape 0 -0.33, +0.33 Both 2p2h shape: NN or Np nucleon pair
@@
RES_E_b_C_numu 25 MeV -24.9, +24.9 CH Resonant SPP $E_{b}$ value (MeV) for muon neutrinos on Carbon
RES_E_b_O_numu 25 MeV -24.9, +24.9 Water Resonant SPP $E_{b}$ value (MeV) for muon neutrinos on Oxygen
RES_E_b_C_numubar 25 MeV -24.9, +24.9 CH Resonant SPP $E_{b}$ value (MeV) for $\overline{\nu}_{\mu}$ on Carbon
RES_E_b_O_numubar 25 MeV -24.9, +24.9 Water Resonant SPP $E_{b}$ value (MeV) for $\overline{\nu}_{\mu}$ on Oxygen
BgSclRES 1.3 -0.15, +0.15 Both I½ non-resonant background
CA5RES 1.01 -0.15, +0.15 Both CA5(0): value at $Q^{2}=0$ of the axial form factor
MaRES 0.95 GeV -0.15, +0.15 Both RES axial mass
@@
PionFSI_AbsProb 1.404 -0.432, +0.432 Both Pion FSI absorption (FEFABS)
PionFSI_CExHighMomProb 1.8 -0.288, +0.288 Both Pion FSI Single charge exchange (high energy) (FEFCXH)
PionFSI_CExLowMomProb 0.697 -0.305, +0.305 Both Pion FSI Single charge exchange (low energy) (FEFCX)
PionFSI_InelProb 1.002 -1.101, +1.101 Both Pion FSI Hadron (N+n pi) production (FEFINEL)
PionFSI_QEHighMomProb 1.824 -0.859, +0.859 Both Pion FSI QE scattering (high energy) (FEFQEH)
PionFSI_QELowMomProb 1.069 -0.313, +0.313 Both Pion FSI QE scattering (low energy) (FEFQE)
TwkDial_FateNucleonFSI 0 -0.3, +0.3 Both Fate of final state nucleons after the FSI interactions
@@
CC_DIS_norm_nu 1 -0.15, +0.15 Both Normalisation of CC DIS for neutrinos
CC_DIS_norm_nubar 1 -0.15, +0.15 Both Normalisation of CC DIS for anti-neutrinos
CC_MultiPi_norm_nu 1 -0.2, +0.2 Both Multi-pion production normalisation (nu)
CC_MultiPi_norm_nubar 1 -0.2, +0.2 Both Multi-pion production normalisation (nubar)
Table 3: List of all systematic parameters from neutrino interaction models that were included in the analysis. The column ‘Dial Name’ is the name of the parameter as it appears in the source code. The column ‘Prior’ is the parameter’s prior value before variation. The column ‘Uncertainty’ represents the $\pm 1\sigma$ prior uncertainty. The column ‘Target’ indicates which type of target material is affected, ”CH” is a shorthand for hydrocarbon, while ”Both” refers to both water and hydrocarbon. The column ‘Description’ contains a short description of the meaning of the parameter. No single parameter dominates the systematic uncertainty from the neutrino interaction models, in the sense that no parameter contributes more than 50% to the total systematic error.